OpenStudy (anonymous):

TRUE OR FALSE: The following is an arithmetic sequence. 2, 4, -9, 5

jimthompson5910 (jim_thompson5910):

Find the difference between each term: 5 - (-9) = ??? -9 - 4 = ??? 4 - 2 = ??? Are these differences the same?

OpenStudy (anonymous):

NO?

jimthompson5910 (jim_thompson5910):

you are correct, 5 - (-9) = 14 while -9 - 4 = -13 So this is NOT an arithmetic sequence.

OpenStudy (anonymous):

thank you!:)

jimthompson5910 (jim_thompson5910):

sure thing